Owl-Eye Radiologists (OR) does various types of diagnostic imaging. Radiologists perform tests using sophisticated equipment. OR's management wants to compute the costs of performing tests for two different types of patients: those who are hospitalized (including those in emergency rooms) and those who are not hospitalized but are referred by physicians. The data for June for the two categories of patients follow.

Hospital Patients Other Patients Total
Units (i.e., procedures) produced 680 830 1,510
Equipment-hours used 240 110 350
Direct labor-hours 500 180 680
Direct labor costs $ 40,000 $ 10,800 $ 50,800
Overhead costs $ 49,730

The accountant first assigns overhead costs to two pools: overhead related to equipment-hours and overhead related to labor-hours. The analysis of overhead accounts by the cost accountant follows.

Account Amount Related to:
Utilities $ 4,850 Equipment-hours
Supplies 12,620 Labor-hours
Indirect labor and supervision 22,060 Labor-hours
Equipment depreciation and maintenance 6,850 Equipment-hours
Miscellaneous 3,350 Equipment-hours


Required:

a. Compute the predetermined overhead rates assuming that Owl-Eye Radiologists uses equipment-hours to allocate equipment-related overhead costs and labor-hours to allocate labor-related overhead costs.

b. Compute the total costs of production and the cost per unit for each of the two types of patients undergoing tests in June.

A. Equipment related (overhead rate) per hour?

Labor-related (overhead rate) per hour?

B. Total Cost for Hospital Patients? Total Cost for Other Patients?

Cost per Unit for Hospital Patients? Cost per Unit for Other Patients?
Answer #1

a) Overhead rate:

Equipment related = 15050/350 = 43 per hour

Labor related = 34680/680 = 51 per hour

b) Total Cost and per unit

Hospital patients Other patients
Direct labor 40000 10800
Overhead related to equipment hour 10320 4730
Overhead related labor hour 25500 9180
Total Cost 75820 24710
Unit 680 830
Cost per unit 111.50 29.77
Answer #2

a) 4,850+ 6,850+ 2,250= 15,050

12,620+ 22,060= 34,680


b) 240*43= 10,320                                  110*43= 4,730         

    500*51= 25,500                                   180*51= 9,180

10,320+25,500= 75820                         4,730+ 9,180= 24,710

75820/680= 111.5                                  24,710/830= 29.7710843373 or 29.77